OpenStudy (anonymous):

what are the steps to factoring this: 8x^4+10x^2-3

OpenStudy (anonymous):

you can use substitution let u=x^2 then factor 8u^2+10u-3 =(2u+3)(4u-1) u=x^2, substitute x^2 back in for u =(2x^2+3)(4x^2-1) you can factor further because the second binomial is a difference of 2 squares =(2x^2+3)(2x-1)(2x+1)